Assume an economy where there are two producers: a wood producer and a chair producer. In a given year, the chair producer makes 25 chairs, of which 5 are sold to the wood producer maker and $100 each, 10 chairs are sold to Tibet and 5 are stored. The wood producer sells 1,000 pounds of wood at $4 each, 100 of which are sold to the chair producer. Compute GDP in this economy using the product approach and the expenditure approach.
Answer #1

GDP using product approach = total value of domestic product manufactured = 20 x 100 (only final goods) + 900 x 4 (only final goods) = 2000 + 3600 = 5600

GDP using expenditure approach = total expenditure (total sales net of intermediate goods) = 25 x 100 - 100 x 4 + 1000 x 4 - 5 x 100 = 5600

In both cases goods sold to each other are intermediary and should be deducted in total value.